# Estimating Spatially Resolved Radiation Fields Using Neural Networks

### v2

@@ 2.4 Training pipeline
To ensure that the models are all comparable, the training pipeline for all models was identical. Regardless, whether the model itself was capable of predicting partial or whole fields, the evaluation of the loss functions and the sub sequential update of the weights was performed in the same way. All models were trained for a maximum of $200$ epochs, which could be prematurely aborted when the validation loss was not improving for the last $10$ epochs. For the training loop, we used the pytorch lightning [37] framework which implements the Cross-Validation policy with a final test using a dedicated validation and test dataset. The ratio of the split into training, validation and test dataset was $70\,\%$, $15\,\%$ and $15\,\%$. At first, a number of fields were loaded from the dataset that are transformed to pytorch tensors, preserving the channels and single layers, and as such they were batched together. We used a physical batch size of $4$ full radiation fields. Multiple batches of those predicted and ground-truth fields were aggregated together to an effective batch size of $64$ for training. We applied data pre-processing to each field prior to the prediction. This pre-processing included not only the option of joining the direct beam and scattered field channels, but also normalizing the ground-truth and input tensors. Finally, the loss functions for each of the neural networks normalized outputs (fluence and spectra) are calculated. The metrics are calculated on the inverse-transformed data in the original data space.
@@
2.4.1 Loss functions.
Across all our models we used the same two loss functions, that we optimized concurrently: $\pazocal{L}_{\Phi^{3}_{\mathrm{\gamma}}}$ the loss regarding the volumetric fluence $\Phi^{3}_{\mathrm{\gamma}}$ and $\pazocal{L}_{p(E_{\mathrm{\gamma}})}$ the loss regarding the local photon spectrum $p(E_{\mathrm{\gamma}})\in\mathbb{R}^{32}$ for $\sum_{i=0}^{31}p(E_{\mathrm{\gamma},i})=1$.
$\pazocal{L}_{\Phi^{3}_{\mathrm{\gamma}}}(t,p)$ $=\frac{1}{3}\Bigl(\mathrm{L_{1}}(t,p)+\mathrm{L_{2}}(t,p)+\bigl(1-\mathrm{SSIM}(t,p)\bigr)\Bigr),\$$\mathrm{with}\ t,p\in\mathbb{R}$
@@
With $t$ being the ground truth (target) and $p$ being the prediction of the neural network. Here, $L_{1}$ and $L_{2}$ refer to the L1- and L2-Norms using mean aggregation to produce a scalar loss for higher dimensional inputs.
2.4.2 Hyperparameter tuning.
@@
feature fusing [Concatenation, FiLM, ResFiLM, GMU]
normalizer $\mathrm{MaxNorm}_{0}^{1}$, $\mathrm{MaxNorm}_{-1}^{+1}$, $\mathrm{MaxLogNorm}_{10^{0}}$, $\mathrm{MaxLogNorm}_{10^{3}}$]
@@ 2.5 Validation Metrics
$\mathrm{Spec}_{acc}=\frac{\sum_{i=1}^{n}{\min(P_{i},T_{i})}}{\sum_{i=1}^{n}{P_{i}+T_{i}-\min(P_{i},T_{i})}}$
@@ 3 Results
Table 3: Reported reconstruction metrics for each network in its optimal configuration of hyperparameters. All metrics were calculated from the test dataset section, grouped by used dataset.
@@
DS-01 SRBF $\mathbf{96.4}\,\%$ $\mathbf{95.5}\,\%$ $\mathbf{0.961}$ $\mathbf{99.84\,\%}$ $\mathbf{99.98}\,\%$ $\mathbf{84.2}\,\%$
@@
SPERF $\mathbf{97.3}\,\%$ $\mathbf{96.2}\,\%$ $\mathbf{0.952}$ $\mathbf{99.7\,\%}$ $\mathbf{99.9}\,\%$ $\mathbf{86.2}\,\%$
@@
In general, the qualities of the developed models by this research work are condensed in Tab. 3. Therein, each metric value belongs to a network, trained, tested and optimized in terms of hyperparameters on one of our datasets. The first dataset DS-01, was used to train our base FCNN architecture called SRBFNet. We compared the results of that network against the implementation of the U-Net that we introduced previously as Beam-to-Scatter (B2S) U-Net. As this network does not need any architectural changes to be applied on multiple datasets, as its inputs are the raw primary beam layer of a radiation field together with the X-ray tube output spectra, we used this as a reference for each dataset and variant of our FCNN architecture. We report the best metric values our networks achieved after we optimized the models hyperparameters by using Optuna as mentioned in the previous section. The specific found hyperparameters for each network, are listed in Tab. 4.
On DS-01, one can easily see, that the FCNN architecture outperforms the U-Net architecture. Even though, the U-Net also learns the structure of the radiation field, the edges between primary beam and scattered radiation are not reproduced as sharply as the FCNN does. This observation can be reproduced for DS-02. Additionally, we intended to highlight the effectiveness of our spectra encoding for the expected air kerma rates $\dot{K}_{\mathrm{air}}$. Thus, we trained the base SRBFNet model, which cannot handle different output spectra of the X-ray tube, on DS-02. The effectivity of our spectra encoding can be seen by comparing SRBFNet against SPERFNet, where we get a nominal difference of $12.9\,\%$ for $\mathrm{SMAPE}_{acc}^{scatter}$ just from the missing spectra information. Furthermore, we can conclude, that training our FCNN architecture on a bigger dataset using more degrees of freedom, namely the X-ray tube spectrum, actually improves the $\mathrm{SMAPE}_{acc}$, even though the spatial resolution got slightly, but not notably, worse.
@@
Dataset Model Model Width L $l_{max}$ $\mathrm{dim}(\mathbf{g}_{\mathrm{spec}})$ Norm Conditioning
DS-01 SRBF 192 10 4 - $\mathrm{MaxNorm}_{0}^{1}$ FiLM
B2S U-Net - - - - $\mathrm{MaxNorm}_{0}^{1}$ -
DS-02 SRBF 192 10 4 - $\mathrm{MaxNorm}_{0}^{1}$ FiLM
SPERF 192 10 4 16 $\mathrm{MaxNorm}_{0}^{1}$ FiLM
B2S U-Net - - - - $\mathrm{MaxNorm}_{0}^{1}$ FiLM
DS-03 PBRF 192 14 4 16 $\mathrm{MaxNorm}_{0}^{1}$ FiLM
B2S U-Net - - - - $\mathrm{MaxNorm}_{0}^{1}$ FiLM
Table 5: Inference durations per field ($50^{3}$ voxels) of each model, regarding a specific feature fusing method, using non-compiled pytorch models and a NVIDIA RTX 3090 GPU. All models were using the following configuration with a model width of $192$ and encoding parameters $l_{max}=4\land\mathrm{L}=10$ for SRBFNet and SPERFNet, with one exception for PBRFNet, which was using L = $14$.
Model Feature fusing Inference duration per field $\mathrm{SMAPE}_{acc}^{scatter}$
SRBF Concat $20.86\,\mathrm{ms}\ \pm\ 2.81\,\mathrm{ms}$ $94.3\,\%$
FiLM $20.55\,\mathrm{ms}\ \pm\ 1.87\,\mathrm{ms}$ $\mathbf{95.5}\,\%$
ResFiLM $24.46\,\mathrm{ms}\ \pm\ 1.82\,\mathrm{ms}$ $94.4\,\%$
GMU $41.00\,\mathrm{ms}\ \pm\ 2.55\,\mathrm{ms}$ $88.7\,\%$
SPERF Concat $21.76\,\mathrm{ms}\ \pm\ 2.57\,\mathrm{ms}$ $95.2\,\%$
FiLM $21.76\,\mathrm{ms}\ \pm\ 2.57\,\mathrm{ms}$ $\mathbf{96.2}\,\%$
ResFiLM $24.71\,\mathrm{ms}\ \pm\ 1.63\,\mathrm{ms}$ $95.4\,\%$
GMU $41.71\,\mathrm{ms}\ \pm\ 2.51\,\mathrm{ms}$ $93.7\,\%$
PBRF Concat $21.03\,\mathrm{ms}\ \pm\ 1.31\,\mathrm{ms}$ $79.2\,\%$
FiLM $22.17\,\mathrm{ms}\ \pm\ 2.54\,\mathrm{ms}$ $\mathbf{84.8}\,\%$
ResFiLM $24.32\,\mathrm{ms}\ \pm\ 1.88\,\mathrm{ms}$ $67.8\,\%$
GMU $40.43\,\mathrm{ms}\ \pm\ 2.42\,\mathrm{ms}$ $40.9\,\%$
@@ 4 Discussion
The configuration of the best performing models during the hyperparameter tuning are listed in Tab. 4. From those results, we can conclude, that for the fluence prediction, the same data range and activation function as used for NeRF-like approaches is beneficial, which is using sigmoid $\sigma$ together with a linear maximum normalization between $[0,1]$. Moreover, when additionally reviewing the table of Tab. 3, we are convinced, that using a model width of $192$ and a location encoding with frequencies within a range of $10$ to at max $14$ is sufficient for the presented use case. Consistently over all datasets, the angular resolution of the beam needed to solve this problem was found to be low enough, that a harmonics count of $4$ for the spherical harmonics encoding of the beam direction is sufficient. Further, two plain FiLM layers for feature fusing, one at an early and one at a late position in the network, was proven to be most effective for this use case compared to using GMU layers, Residual-FiLM layers and feature concatenation. Even though, the impact of the FiLM layers on the inference duration is measurable, it is small compared to the concatenation that is commonly used in the field of Novel View Synthesis, as it can be seen in Tab. 5. In that table, we concentrate on the differences regarding $\mathrm{SMAPE}_{acc}^{scatter}$, as this metric is reflecting the accuracy on the most interesting part of the predicted volume for the computational dosimetry. That is because the staff, during an IR procedure, will be mainly monitored in the $\mathrm{SMAPE}_{acc}^{scatter}$ volume. From Tab. 5, one can also draw the conclusion, that the use of FiLM layers did not only achieve the best accuracies, but also the shortest inference durations compared to GMU or Residual-FiLM. Moreover, we demonstrated, that the radiation transport simulation can be learned by the same deep, but narrow structure that is used to learn light transfer as in the use case of NVS. As the provided implementation does not use custom, native GPU codes like implementations in CUDA or OpenCL, its inference time does not enable real-time usage in the sense of the required $8\,\mathrm{ms}$ to $11\,\mathrm{ms}$ for VR- or AR-applications. Nevertheless, an inference time of about $20\,\mathrm{ms}$ is suitable for interactive applications. After all, with the provided FCNN models, state-of-the-art inferences times, that are suitable for the VR- and AR-application, should be achievable by implementing the network as CUDA or OpenCL kernels in the future, as has already been demonstrated by InstantNGP.
@@ 5 Conclusion
We have presented three datasets for training machine learning and especially deep learning algorithms on spatially resolved and physically accurate radiation field distributions as they could arise during IR procedures. Additionally, we have implemented and compared different variants of fully connected and convolutional neural networks, following the architectural design of U-Nets and NeRF-like architectures. Thereby, we have proven NeRF-like architectures to be efficient and capable of being used for radiation protection purposes and training of medical staff in the context of IR procedures. We have also pointed out, which design decisions are effective in this special context. Our presented datasets together with the presented findings on the architectural decisions for neural reconstruction networks create a profound foundation for enabling further research on this topic, e.g., extending our or similar approaches by uncertainty handling in neural networks or incorporation of dynamic scatter geometry.

### v3

@@ 2.3 Training pipeline
To ensure that the models are all comparable, the training pipeline for all models was identical. Regardless, whether the model itself was capable of predicting partial or whole fields, the evaluation of the loss functions and the sub sequential update of the weights was performed in the same way. All models were trained for a maximum of $200$ epochs, which could be prematurely aborted when the validation loss was not improving for the last $10$ epochs. For the training loop, we used the pytorch lightning [8] framework in version 2.5.6 which implements the Cross-Validation policy with a final test using a dedicated validation and test dataset. The ratio of the split into training, validation and test dataset was $70\,\%$, $15\,\%$ and $15\,\%$. At first, a number of fields were loaded from the dataset that are transformed to pytorch tensors, preserving the channels and single layers, and as such they were batched together. We used a physical batch size of $4$ full radiation fields. Multiple batches of those predicted and ground-truth fields were aggregated together to an effective batch size of $64$ for training. We applied data pre-processing to each field prior to the prediction. This pre-processing included not only the option of joining the direct beam and scattered field channels, but also normalizing the ground-truth and input tensors. Finally, the loss functions for each of the neural networks normalized outputs (fluence and spectra) are calculated. The metrics are calculated on the inverse-transformed data in the original data space.
@@
2.3.1 Loss functions
Across all our models we used the same two loss functions, that we optimized concurrently: $\pazocal{L}_{\Phi^{3}_{\mathrm{\gamma}}\penalty 10000\ }$ the loss regarding the volumetric fluence $\Phi^{3}_{\mathrm{\gamma}}$ and $\pazocal{L}_{p(E_{\mathrm{\gamma}})}$ the loss regarding the local photon spectrum $p(E_{\mathrm{\gamma}})\in\mathbb{R}^{32}$ for $\sum_{i=0}^{31}p(E_{\mathrm{\gamma},i})=1$.
$\pazocal{L}_{\Phi^{3}_{\mathrm{\gamma}}\penalty 10000\ }(t,p)$ $=\frac{1}{3}\Bigl(\mathrm{L_{1}}(t,p)+\mathrm{L_{2}}(t,p)+\bigl(1-\mathrm{SSIM}(t,p)\bigr)\Bigr),\$$\mathrm{with}\ t,p\in\mathbb{R}$
@@ 2.3.1 Loss functions
With $t$ being the ground truth (target) and $p$ being the prediction of the neural network. A quick exploration of the parameter space had lead to the presented ratio between Wasserstein and $L_{1}$-Loss and was beneficial for achieving satisfactory results. Here, $L_{1}$ and $L_{2}$ refer to the L1- and L2-Norms using mean aggregation to produce a scalar loss for higher dimensional inputs.
2.3.2 Hyperparameter tuning
@@
feature fusion [Concatenation, FiLM, ResFiLM, GMU]
@@ 2.4 Validation Metrics
$\mathrm{SSIM}(p,t)=\frac{(2\mu_{p}\mu_{t}+C_{1})(2\sigma_{pt}+C_{2})}{(\mu_{p}^{2}+\mu_{t}^{2}+C1)(\sigma_{p}^{2}+\sigma_{t}^{2}+C_{2})}$
with $\mu_{x}$ being the mean of the window $x$, $\sigma_{x}$ being the variance of the window $x$, and $\sigma_{xy}$ being the covariance matrix of the two windows $x$ and $y$. The $\mathrm{SSIM}(p,t)$ is calculated for small windows $p$ and $t$ of the compared volumes $\mathbf{P}$ and $\mathbf{T}$, and the global SSIM is calculated by the weighted sum of all windows. $C_{1}$ and $C_{2}$ are calculated from the dynamic range of the compared volumes as stabilization constants to prevent terms from degrading to zero.
@@
$\mathrm{Spec}_{acc}=\frac{\sum_{i=1}^{n}{\min(P_{i},T_{i})}}{\sum_{i=1}^{n}{\bigl(P_{i}+T_{i}-\min(P_{i},T_{i})\bigr)}}$
@@ 3 Results
In general, the qualities of the developed models by this research work are condensed in Tab. 4. Therein, each metric value belongs to a network, trained, tested and optimized in terms of hyperparameters on one of our datasets. The first dataset DS-01, was used to train our base FCNN architecture called SRBFNet. We compared the results of that network against the implementation of the U-Net that we introduced previously as Beam-to-Scatter (B2S) U-Net. As this network does not need any architectural changes to be applied on multiple datasets, as its inputs are the raw primary beam layer of a radiation field together with the X-ray tube output spectra, we used this as a reference for each dataset and variant of our FCNN architecture. We report the best metric values our networks achieved after we optimized the models hyperparameters by using Optuna as mentioned in the previous section. The specific found hyperparameters for each network are listed in Tab. 5.
Table 4: Reported reconstruction metrics for each network in its optimal configuration of hyperparameters. All metrics were calculated from the test dataset section, grouped by used dataset. For the training of the models a NVIDIA H100 GPU was used.
@@
DS-01 SRBF $\mathbf{96.4}\,\%$ $\mathbf{95.5}\,\%$ $\mathbf{0.961}$ $\mathbf{99.8\,\%}$ $\mathbf{99.98}\,\%$ $\mathbf{84.2}\,\%$
@@
SPERF $\mathbf{97.1}\,\%$ $\mathbf{96.2}\,\%$ $\mathbf{0.952}$ $\mathbf{99.7\,\%}$ $\mathbf{99.9}\,\%$ $\mathbf{86.2}\,\%$
@@
On DS-01, one can easily see, that the FCNN architecture outperforms the U-Net architecture. Even though, the U-Net also learns the structure of the radiation field, as reflected by the SSIM scores in Tab. 4, the edges between primary beam and scattered radiation are not reproduced as sharply as the FCNN does. This can be concluded from the relation between SMAPE accuracy on both regions and the SSIM scores, which drops less than the accuracy. This observation can be reproduced for DS-02. Additionally, we intended to highlight the effectiveness of our spectra encoding for the expected air kerma rates $\dot{K}_{\mathrm{air}}$. Thus, we trained the base SRBFNet model, which cannot handle different output spectra of the X-ray tube, on DS-02. The effectivity of our spectra encoding can be seen by comparing SRBFNet against SPERFNet, where we get a nominal difference of $12.9\,\%$ for $\mathrm{SMAPE}_{acc}^{scatter}$ just from the missing spectra information. Furthermore, when comparing SRBF on DS-01 with SPERF on DS-02 in Tab. 4, we can conclude, that training our FCNN architecture on a bigger dataset using more degrees of freedom, namely the X-ray tube spectrum, actually improves the $\mathrm{SMAPE}_{acc}$, even though the spatial resolution , according to the drop in the SSIM score, got slightly, but not notably, worse.
@@
From Tab. 6, we can see, that the use of GMU feature fusion is doubling the inference durations and decreases the learning capabilities consistently over all datasets and FCNN based models. The use of FiLM feature fusion, on the other hand, is faster and often consistently beneficial for the SMAPE accuracy metric over the scattered radiation field region. In general, an inference duration for a whole radiation field estimation of about $20\,\mathrm{ms}$ was achieved by all our FCNN based models, when used with concatenation or FiLM based feature fusion. Please note, that the corresponding B2S-U-Net implementation achieved a mean inference duration of $7.5\,\mathrm{ms}\pm 2.0\,\mathrm{ms}$ on the same datasets and hardware using FiLM conditioning in the center layer.
@@
Dataset Model Model Width L $l_{max}$ $\mathrm{dim}(\mathbf{g}_{\mathrm{spec}})$ Conditioning
DS-01 SRBF 192 10 4 - FiLM
B2S U-Net - - - - -
DS-02 SRBF 192 10 4 - FiLM
SPERF 192 10 4 16 FiLM
B2S U-Net - - - - FiLM
DS-03 PBRF 192 14 4 16 FiLM
B2S U-Net - - - - FiLM
Table 6: Inference durations per field ($50^{3}$ voxels) of each model, regarding a specific feature fusion method, using non-compiled pytorch models and a NVIDIA RTX 4090 GPU in a Intel i9-13900K system with $128\,\mathrm{GB}$ of RAM installed. All models were using the following configuration with a model width of $192$ and encoding parameters $l_{max}=4\land\mathrm{L}=10$ for SRBFNet and SPERFNet, with one exception for PBRFNet, which was using L = $14$.
Model Feature fusion Inference duration per field $\mathrm{SMAPE}_{acc}^{scatter}$
SRBF Concat $20.7\,\mathrm{ms}\ \pm\ 2.8\,\mathrm{ms}$ $94.3\,\%$
FiLM $20.6\,\mathrm{ms}\ \pm\ 1.9\,\mathrm{ms}$ $\mathbf{95.5}\,\%$
ResFiLM $24.5\,\mathrm{ms}\ \pm\ 1.8\,\mathrm{ms}$ $94.4\,\%$
GMU $41.0\,\mathrm{ms}\ \pm\ 2.6\,\mathrm{ms}$ $88.7\,\%$
SPERF Concat $21.8\,\mathrm{ms}\ \pm\ 2.6\,\mathrm{ms}$ $95.2\,\%$
FiLM $21.8\,\mathrm{ms}\ \pm\ 2.6\,\mathrm{ms}$ $\mathbf{96.2}\,\%$
ResFiLM $24.7\,\mathrm{ms}\ \pm\ 1.6\,\mathrm{ms}$ $95.4\,\%$
GMU $41.7\,\mathrm{ms}\ \pm\ 2.5\,\mathrm{ms}$ $93.7\,\%$
PBRF Concat $21.0\,\mathrm{ms}\ \pm\ 1.3\,\mathrm{ms}$ $79.2\,\%$
FiLM $22.2\,\mathrm{ms}\ \pm\ 2.5\,\mathrm{ms}$ $\mathbf{84.8}\,\%$
ResFiLM $24.3\,\mathrm{ms}\ \pm\ 1.9\,\mathrm{ms}$ $67.8\,\%$
GMU $40.4\,\mathrm{ms}\ \pm\ 2.4\,\mathrm{ms}$ $40.9\,\%$
@@ 4 Discussion
The configuration of the best performing models during the hyperparameter tuning are listed in Tab. 5. When additionally reviewing the table of Tab. 4, we are convinced, that using a model width of $192$ and a location encoding with frequencies within a range of $10$ to at max $14$ is sufficient for the presented use case. Consistently over all datasets, the angular resolution of the beam needed to solve this problem was found to be low enough, that a harmonics count of $4$ for the spherical harmonics encoding of the beam direction is sufficient. Further, two plain FiLM layers for feature fusion, one at an early and one at a late position in the network, was proven to be most effective for this use case compared to using GMU layers, Residual-FiLM layers and feature concatenation. Even though, the impact of the FiLM layers on the inference duration is measurable, it is small compared to the concatenation that is commonly used in the field of Novel View Synthesis, as it can be seen in Tab. 6. In that table, we concentrate on the differences regarding $\mathrm{SMAPE}_{acc}^{scatter}$, as this metric is reflecting the accuracy on the most interesting part of the predicted volume for the computational dosimetry. That is because the staff, during an IR procedure, will be mainly monitored in the $\mathrm{SMAPE}_{acc}^{scatter}$ volume. From Tab. 6, one can also draw the conclusion, that the use of FiLM layers did not only achieve the best accuracies, but also the shortest inference durations compared to GMU or Residual-FiLM. Moreover, we demonstrated, that the radiation transport simulation can be learned by the same deep, but narrow structure that is used to learn light transfer as in the use case of NVS. As the provided implementation does not use custom, native GPU codes like implementations in CUDA or OpenCL, its inference time does not enable real-time usage in the sense of the required $8\,\mathrm{ms}$ to $11\,\mathrm{ms}$ for VR- or AR-applications. Nevertheless, an inference time of about $20\,\mathrm{ms}$ is suitable for interactive applications. After all, with the provided FCNN models, state-of-the-art inferences times, that are suitable for the VR- and AR-application, should be achievable by implementing the network as CUDA or OpenCL kernels in the future, as has already been demonstrated by InstantNGP [19].
@@ 5 Conclusion
We have presented three datasets for training machine learning and especially deep learning algorithms on spatially resolved and physically accurate radiation field distributions as they could arise during IR procedures. Additionally, we have implemented and compared different variants of fully connected and convolutional neural networks, following the architectural design of U-Nets and NeRF-like architectures. Our results indicate that NeRF-like architectures are effective and suitable for radiation protection applications, and are therefore likely applicable to the training of medical staff in the context of IR procedures. We have also pointed out, which design decisions are effective in this special context. Our presented datasets together with the presented findings on the architectural decisions for neural reconstruction networks create a profound foundation for enabling further research on this topic, e.g., extending our or similar approaches by uncertainty handling in neural networks or incorporation of dynamic scatter geometry.
6 Future work
We leave it as future work to implement the presented FCNN using fused CUDA kernels to improve their performance. In doing so, it is expected that the networks will achieve inference times suitable for AR/VR rendering, as we have presented lightweight NeRF-like architectures with two FiLM layers. We demonstrated that these FiLM layers only add minimal overhead. In general, this approach was already demonstrated for NeRF variants like InstantNGP [19]. Further, we empathize the importance of public available datasets for training and comparing neural networks for dosimetry. Especially, more datasets with increasing realism are needed by adding beam collimation that is independent from the tubes distance from the isocenter or by adding various patient shapes and poses to the dataset generation process.